OpenStudy (anonymous):

Find domain of the following function. write answer in interval notation. y=(x) / (cos x) step by step would be much appreciated

OpenStudy (anonymous):

the only thing you have to worry about is dividing by zero. so the question comes down to "for which x is \[\cos(x)=0\]

OpenStudy (anonymous):

and the answer is, \[\frac{\pi}{2},\frac{3\pi}{2},\frac{5\pi}{2},...\]

OpenStudy (anonymous):

if you want a formula for an odd number times pi over 2 you cana write \[\frac{(2k+1)\pi}{2}\] for any integer k

OpenStudy (anonymous):

because cosx cannot equal to zero? the answer you have has dots at the end,,,,,so it never ends? so its all real numbers besides those?

OpenStudy (anonymous):

yes all real number besides the ones i wrote cosine is periodic and has an infinite number of zeros

OpenStudy (anonymous):

thank you!